OpenStudy (anonymous):

6.03] Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10 3

Directrix (directrix):

Look at the two equations. What would you get if you added the like terms? 2x + 5y = −3 −15x − 5y = −10

OpenStudy (marissalovescats):

I would use elimination because of the like terms :')